The function s(t)=-16t^2+64t+200 gives the height above ground s. in feet of a model rocket launched with a velocity of 64 ft/s from a tower that is 200 feet high. Determine when the rocket hits the ground.
Answer by dabanfield(803)

The function s(t)=-16t^2+64t+200 gives the height above ground s. in feet of a model rocket launched with a velocity of 64 ft/s from a tower that is 200 feet high. Determine when the rocket hits the ground.
We need to find the values of t that make s(t) = 0.
-16t^2 + 64t + 200 = 0
Divide both sides by 8 to simplify the equation:
-2t^2 + 8t + 25 = 0
This is a quadratic equation with a=-2, b=8 and c = 25. Use the quadratic formula to find the positive value of t.